Please help, my mom, dad and brother are no help!
8 over w-10 = 16 over w-5 = z+8 over w

Answer by oberobic(2304)

8 / (w-10) = 16 / ( w-5) = (z+8) / w
.
We know that if a = b and b = c, then a = c.
So you can do this problem in pieces.
.
8 / (w-10) = 16 / ( w-5)
.
cross multiply
.
8(w-5) = 16(w-10)
8w -40 = 16w -160
-8w = -120
.
w = 15
.
16 / ( w-5) = (z+8) / w
.
cross multiply
.
16w = (z+8)(w-5)
.
substitute w=15
.
16(15) = (z+8)(15-5)
.
240 = (z+8)*10
.
240 = 10z +80
10z = 240 - 80
.
z = 16
.
Now that we've found w & z, we can check to see if the three original equations are satisfied.
.
8 / (w-10) = 8/(15-10) = 8/5
.
16 / ( w-5) = 16 /(15-5) = 16/10 = 8/5
.
(z+8) / w = (16+8) / 15 = 24/15 = 8/5
.
Correct.
.
Answer: w=15 and z=16.
.
Done.